# PeersimGym: An Environment for Solving the Task Offloading Problem with Reinforcement Learning

### v1

@@ 5.1 Experimental Setup
The experiments are grounded in four distinct network topologies generated using Ether, each featuring an incremental number of Array of Things (AoT) clusters. These scenarios leverage the realistic topology and workload generation methodologies detailed in Section 4.3. Each AoT cluster, predominantly composed of Single Board Computers (SBCs), includes pairs of SBCs and a base station equipped with an Intel NUC and two GPU units, alongside a remote, more potent server.
@@
The topologies vary in cluster numbers, spanning one to four clusters, and correspondingly in node counts, from 12 to 31. The controller protocol is present in varying quantities (8 to 22 nodes across different topologies). Despite uniform CPU frequencies, the computational power varies with the core count, with the remote server having the highest capacity. The communication capabilities and bandwidths are homogeneous across all nodes. Task generation at each SBC node follows a $Poisson(\lambda)$ distribution, over a simulation episode of $1000$ time steps. Each agent makes an offloading decision in each time step, through $300$ training and $100$ inference episodes. See Table 1 for exact values

### v3

@@ 5.1 Experimental Setup
The experiments are grounded in four distinct network topologies generated using Ether, each featuring an incremental number of Array of Things (AoT) clusters. These scenarios leverage the realistic topology and workload generation methodologies detailed in Sec. 4.3. Each AoT cluster, predominantly composed of Single Board Computers (SBCs), includes pairs of SBCs and a base station equipped with an Intel NUC and two GPU units, alongside a remote, more potent server. The simulation parameters for all scenarios can be found in the Agent Repository.
@@
The topologies vary in cluster numbers, spanning one to four clusters, and correspondingly in node counts, from 12 to 31. The controller protocol is present in varying quantities (8 to 22 nodes across different topologies). Despite uniform CPU frequencies, the computational power varies with the core count, with the remote server having the highest capacity. The communication capabilities and bandwidths are homogeneous across all nodes. Task generation at each SBC node follows a $Poisson(\lambda)$ distribution, over a simulation episode of $1000$ time steps. Each agent makes an offloading decision in each time step, through $300$ training and $100$ inference episodes. See Tab. 1 for exact values  [^3] We provide the in-depth configurations for the environment in the agent repository
@@ 6 Conclusion and Future Work
Acknowledgements This work was partially supported by NOVA LINCS (UI DB/04516/2020) with the financial support of FCT I.P. and Project "Artificial Intelligence Fights Space Debris" No C626449889-0046305 co-funded by Recovery and Resilience Plan and NextGeneration EU Funds, www.recuperarportugal. gov.pt. And, by the European Union (TARDIS, 101093006). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union. Neither the European Union nor the granting authority can be held responsible for them.